OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 4, -14, and 5 + 8i f(x) = x4 - 362.5x2 + 1450x - 4984 f(x) = x4 - 9x3 + 32x2 - 725x + 4984 f(x) = x4 - 67x2 + 1450x - 4984 f(x) = x4 - 9x3 - 32x2 + 725x - 4984

OpenStudy (misty1212):

ok HI!! lets knock this out quick

OpenStudy (anonymous):

Thank you!

OpenStudy (misty1212):

gonna start with \[(x-4)(x+14)\] and then multiply it by a polynomial with zeros at \(5+8i\) and \(5-8i\)

OpenStudy (misty1212):

there are a couple ways to find that quadratic one is hard, one is easy and one is real real easy you pick

OpenStudy (anonymous):

Let's do the real real easy one haha

OpenStudy (misty1212):

ok the quadratic with zeros as \(a+bi\) is \[x^2-2ax+(a^2+b^2)\] in your case \(a+bi=5+8i\) so the quadratic is \[x^2-2\times 5x+(5^2+8^2)=x^2-10x+89\]

OpenStudy (misty1212):

final job is to multiply \[(x-4)(x+14)(x^2-10x+89)\] the algebra is a real pain so lets let the computer do it

OpenStudy (anonymous):

Okay wait

OpenStudy (anonymous):

x^4−67x^2+1450x−4984?

OpenStudy (misty1212):

looks good to me

OpenStudy (misty1212):

easy enough right?q

OpenStudy (misty1212):

cute dog btw

OpenStudy (anonymous):

Yes thank you so much!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]